# FIRE: A Failure-Adaptive Reinforcement Learning Framework for Edge Computing Migrations

### v2

@@ 6.3 Proofs of Boundedness and Convergence to Optimality
Algorithm 1’s update equation is
$\begin{split}\hat{Q}^{t+1}(s^{t},a^{t})&\leftarrow\hat{Q}^{t}(s^{t},a^{t})+\alpha[w_{t}(r^{t+1}+\gamma\hat{Q}^{t}(s^{t+1},a^{*}))\\ &-\hat{Q}^{t}(s^{t},a^{t})],\end{split}$ (40)
where $a^{*}=\text{argmax}_{b}\hat{Q}(s^{t+1},b)$.
By definition, the optimal policy $Q^{*}$ of the true system with transition probability $p(s^{\prime}|s,a)$ is a fixed point of the Bellman optimality equation:
$Q^{*}(s,a)=\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\max_{b}Q^{*}(s^{\prime},b)].$
$\hat{p}(s^{\prime}|s,a)$ is the transition probability in our simulator which incorporates importance sampling. $\hat{p}(s^{\prime}|s,a)=\begin{cases}(1-\hat{\epsilon}(s))h(\tilde{s}|s,a),&\text{if}\ s^{\prime}\notin T\\ \hat{\epsilon}(s)h(\tilde{s}|s,a),&\text{if}\ s^{\prime}\in T,\end{cases}$ as importance sampling of rare events takes place according to $\hat{\epsilon}(s)$. Without loss of generality, when we look at the case where $s^{\prime}$ is a rare state ($s^{\prime}\in T$), we have
$\hat{p}(s^{\prime}|s,a)w_{t}(s)$ $=\hat{\epsilon}(s)h(\tilde{s}|s,a)\frac{\epsilon(s)}{\hat{\epsilon}(s)}$
$=p(s^{\prime}|s,a),$
the actual transition probability of the system. Therefore, since $p(s^{\prime}|s,a)=\hat{p}(s^{\prime}|s,a)w_{t}(s)$, the optimal policy $Q^{*}$ is a fixed point of the following equation with our simulator’s transition probability $\hat{p}(s^{\prime}|s,a)$ and the importance sampling correction weight $w_{t}(s)$:
$Q^{*}(s,a)=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}Q^{*}(s^{\prime},b)].$ (41)
We define the operator $\mathbf{H}$ for a generic function $\hat{Q}_{i}$ as
$(\mathbf{H}\hat{Q})(s,a)=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}(s^{\prime},b)].$ (42)
We have shown above that $\mathbf{H}Q^{*}=Q^{*}$. Now we show that $\mathbf{H}$ is a contraction mapping. Note that both $\hat{p}(s^{\prime}|s,a)$ and $w_{t}(s)$ are functions of $\hat{Q}$, because they are functions of $\hat{\epsilon}$, which is obtained as a function of $\hat{Q}$ (algorithm lines 10-15). Nevertheless, we are able to prove contraction mapping because the product $\hat{p}(s^{\prime}|s,a)(\hat{Q})\times w_{t}(s(\hat{Q}))$ is the true transition probability $p(s^{\prime}|s,a)$.
$||\mathbf{H}\hat{Q}_{1}-\mathbf{H}\hat{Q}_{2}||_{\infty}$ (43)
$=\max_{s,a}|\sum_{s^{\prime}}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}_{1}(s^{\prime},b)]$
$-\sum_{s^{\prime}}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}_{2}(s^{\prime},b)]|$
$\leq\max_{i,a}\gamma\sum_{s^{\prime}\in S}p(s^{\prime}|i,a)|\max_{y}\hat{Q}_{1}(s^{\prime},y)-\max_{y}\hat{Q}_{2}(s^{\prime},y)|$
$\leq\max_{i,a}\gamma\sum_{s^{\prime}\in S}p(s^{\prime}|i,a)|\max_{y}|\hat{Q}_{1}(s^{\prime},y)-\hat{Q}_{2}(s^{\prime},y)||$
$\leq\max_{i,a}\gamma\sum_{s^{\prime}\in S}p(s^{\prime}|i,a)\max_{z,y}|\hat{Q}_{1}(z,y)-\hat{Q}_{2}(z,y)|$
$=\max_{i,a}\gamma\sum_{s^{\prime}\in S}p(s^{\prime}|i,a)||\hat{Q}_{1}-\hat{Q}_{2}||_{\infty}$
$=\gamma||\hat{Q}_{1}-\hat{Q}_{2}||_{\infty},$ (44)
where $\gamma\in[0,1)$ according to the theorem conditions. The first inequality is true by the triangle inequality. The second inequality holds because $\max_{x}f(x)-\max_{x}g(x)\leq\max_{x}|f(x)-g(x)|$. Hence, we have shown that operator $H$ is a contraction mapping.
To prove convergence of our algorithm, we use the result from Corollary 3. Firstly, we subtract $Q^{*}(s^{t},a^{t})$ from both sides of Eq. (40), and letting $\triangle_{t}(s,a)=\hat{Q}^{t}(s,a)-Q^{*}(s,a)$, we have:
$\triangle_{t+1}(s^{t},a^{t})$ $=(1-\alpha^{t})\triangle_{t}(s^{t},a^{t})+\alpha^{t}[w^{t}(r^{t+1}$
$+\gamma\max_{b}\hat{Q}^{t}(s^{t+1},b))-Q^{*}(s^{t},a^{t})],$
where $r^{t+1}=r(s^{t},a^{t},s^{t+1})$. Let $G_{t}(s,a)=w^{t}[r(s,a,X(s,a))+\gamma\max_{b}\hat{Q}^{t}(X(s,a),b)]-Q^{*}(s,a),$ where $X(s,a)$ is a random sample obtained from the markov chain with state space $S$ and transition probability $\hat{p}(s^{\prime}|s,a)$ (the importance sampling modified transition probability). Taking the expectation, we have
$\mathbb{E}[G_{t}(s,a)]$ $=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w^{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}^{t}(s^{\prime},b)]$
$-Q^{*}(s,a)$
$=\mathbf{H}\hat{Q}^{t}(s,a)-Q^{*}(s,a)$
$=\mathbf{H}\hat{Q}^{t}(s,a)-\mathbf{H}Q^{*}(s,a).$
The last equality holds because $Q^{*}$ is a fixed point of operator $\mathbf{H}$, as we have shown in Eqs. (41) and (42). Since we have shown that $\mathbf{H}$ is a contraction mapping above, we have
$||\mathbb{E}[G_{t}(s,a)]||_{\infty}\leq\gamma||\hat{Q}^{t}(s,a)-Q^{*}||_{\infty}=\gamma||\triangle_{t}||_{\infty}.$ (45)
It remains to show that $\textbf{var}[G_{t}(s,a)]\leq C(1+||\triangle_{t}||^{2}),\text{for}\ C>0$, where var refers to the variance.
$\textbf{var}[G_{t}(s,a)]$
$=\mathbb{E}[\Bigl{(}w^{t}[r(s,a,X(s,a))+\gamma\max_{b}\hat{Q}^{t}(X(s,a),b)]$
$-\mathbf{H}\hat{Q}^{t}(s,a)\Bigr{)}^{2}]$
$=\textbf{var}\Bigl{(}w^{t}[r(s,a,X(s,a))+\gamma\max_{b}\hat{Q}^{t}(X(s,a),b)]\Bigr{)}$
Since $w^{t}$ and $r$ are bounded, we have
$\mathbf{var}[G_{t}(s,a)]\leq C(1+||\triangle_{t}||^{2}_{\infty})$
for some constant $C>0$. With this, we have shown the convergence of Algorithm ImRE. ∎
@@
Rare Event Adaptive Eligibility Traces Algorithm (ETAA, Algorithm 2): To speed up convergence for the tabular Q learning algorithm, we propose another variant of Algorithm 1 using eligibility traces. This algorithm (ETAA) is a combination of our proposed importance sampling based Q-learning Algorithm 1 and Watkin’s $Q(\lambda)$ [48], which is an algorithm which combines Q-learning with eligibility traces. In this algorithm, $\hat{Q}(s,a)$ is updated for many combinations of $(s,a)$, at every time-slot, which helps to speed up convergence. Firstly, when we visit state $s^{t}$ and take action $a^{t}$, we add one to the eligibility traces, as seen in line 5. Next, the eligibility traces will be decayed according to the factor $\lambda$:
$e(s,a)\leftarrow\gamma\lambda e(s,a)w_{t},$ (46)
as seen in line 14, where $\gamma$ is the reward discount factor and $w_{t}$ is the importance weight. This means that the more recently we have visited $(s,a)$, the larger $e(s,t)$ will be, as less decaying has taken place. Next, the larger $e(s,t)$ is, the more $\hat{Q}(s,a)$ is updated by the TD-error $\triangle_{t}$, as seen in line 7. If the action taken is not the greedy action, i.e., exploration has taken place, the eligibility traces will be set to 00, as seen in line 16. The rest of the algorithm follows Algorithm 1.
Algorithm 2 ETAA: Eligibility Traces Rare Events-Adaptive Algorithm
1:Initialise: Same variables as Algorithm 1, along with eligibility traces $e(s,a)$.
2:Select the initial state $s^{0}$ and action $a^{0}$.
3:for all timeslots $t$ do
4:     The procedure follows that of Algorithm 1.
5:     When we visit $s^{t}$ and take action $a^{t}$, $e(s^{t},a^{t})\leftarrow e(s^{t},a^{t})+1.$
6:     $a^{t+1}\leftarrow\beta-\text{greedy}(s^{t+1}).$
7:     $a^{*}\leftarrow\text{argmax}_{b}\hat{Q}(s^{t+1},b).$
8:     if  $\hat{Q}(s^{t+1},a^{*})=\hat{Q}(s^{t+1},a^{t+1})$ then
9:         $a^{*}\leftarrow a^{t+1}$
10:     end if
11:     for all $(s,a)$ do
12:         $\hat{Q}(s,a)\leftarrow\hat{Q}(s,a)+\alpha e(s,t)\triangle_{t}$.
13:         if $a^{t+1}=a^{*}$ then
14:              $e(s,a)\leftarrow\gamma\lambda e(s,a)w_{t}$
15:         else
16:              $e(s,a)\leftarrow 0.$
17:         end if
18:     end for
19:end for
Corollary 5 states that ETAA remains bounded, for the case where $\lambda=1$. Its proof is in our technical report [45].
Corollary 5.
The eligibility traces Q-Learning algorithm (ETAA), Algorithm 2 is bounded for the case $\lambda=1$.
7 Algorithms for large state and action spaces
@@
To handle large state spaces, as might be found in real-world edge computing networks that span one or more cities and surrounding suburbs, we propose two variants of Algorithm 1 using deep Q-learning (Algorithm ImDQL) and actor critic (Algorithm ImACRE) respectively. The codes are available online at [49].
@@
Rare Event Adaptive Deep Q-learning algorithm (ImDQL, Algorithm 3): We modify and combine the classic deep reinforcement learning algorithm in [50] with our importance sampling based rare events adaptive algorithm (Algorithm 1) by proposing an importance sampling based loss function.
@@
Algorithm 3 ImDQL: Importance Sampling for Rare Events using Deep Q Learning
@@
$L(\boldsymbol{\theta})=\mathbb{E}_{s,a,s^{\prime}}[w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))-Q(s,a,\boldsymbol{\theta})]^{2},$ (47)
where $\boldsymbol{\theta}^{-}$ are the weights of the target neural network, $\boldsymbol{\theta}$ are the weights of the predicted neural network, which are updated every iteration, and $w_{t}$ is the importance sampling weight. The loss function is optimized via gradient descent in line 13. Here, we are trying to minimize the difference between the target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ and the prediction $Q(s,a,\boldsymbol{\theta})$. Because the target variable does not change in deep learning, here the target network is updated only every $K$ iterations (line 16), while the prediction network is updated every iteration. Having the target fixed for awhile helps in having stable training. The key idea which differentiates our algorithm from traditional deep reinforcement learning algorithms is that our target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ is “error corrected” by the importance sampling weight $w_{t}$, and the rare event distribution is updated via $T,U$ and $\hat{\epsilon}$. The variables $T$, $U$, $\hat{\epsilon}$ and $w_{t}$ are updated in the same manner as the tabular algorithm (Algorithm 1), through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t+1}$. The actions are also selected in the same manner as Algorithm 1, through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t}$.
Algorithm 4 ImACRE: Importance Sampling based Advantage Actor Critic for Rare Events
@@
Rare Event Adaptive Actor Critic Algorithm (ImACRE, Algorithm 4): Neural networks are used to approximate the “actor” and “critic” functions [51]. The critic function estimates the value (network $c$), based on which the actor function updates the policy (network $\mathbf{\theta}$). The advantage function $A(s^{t},a^{t})$ (line 8) measures how good taking action $a^{t}$ is, compared to taking the average action, at state $s^{t}$. It is a function of the reward $r^{t+1}$, the values $V(s^{t})$ and $V(s^{t+1})$ (obtained from the critic network $\boldsymbol{c}$ , line 7). It is corrected by the importance sampling correction weight $w^{t}(s^{t},a^{t},s^{t+1})$ (line 8), because we perform importance sampling of rare events. Each iteration, $A(s^{t},a^{t})$ will be stored in the buffer (line 9). The variables $T,U,\hat{\epsilon}$ and $w_{t}$ are updated in the same manner as the tabular algorithm (Algorithm 1). At the end of an epoch ($K$ timeslots), both the actor and critic networks are optimized, via gradient descent. The gradients for the actor and critic networks, respectively, are:
@@
8 Adapting to Heterogeneous Risk Tolerances
Edge computing users, application vendors, and network providers have different tolerances towards server failures and the resulting higher latency which occur. Some applications are highly latency sensitive, and any server failure has a significant impact on safety and smooth functioning of the application, e.g., obstacle detection in autonomous vehicles, or immersive outdoor sport applications. Other users with less urgent jobs may see less of a need for a failure-aware backup placement resilient solution, due to the extra costs when storing and migrating backups. To deal with these heterogeneous risk tolerances, one potential solution is to alter the failure cost when training the algorithm. Nevertheless, the same rare event with the same objective cost to the system may inspire different risk tolerances and different willingness to incur a preparation cost amongst different users. Therefore, we propose a mechanism that can cater to edge computing users of heterogeneous risk tolerances. It takes the users’ risk tolerance towards server failures as an input, and derives a joint service placement and backup placement policy, based on the users’ risk tolerance. With RiTA (Risk Tolerance-Adaptive Algorithm), we also do not need to re-train a new policy $Q_{riskAware}$ for every possible level of risk, which scales better to a wider variety of users.
@@ 8 Adapting to Heterogeneous Risk Tolerances
We present this algorithm in Algorithm 5 (RiTA). Firstly, RiTA takes the user’s risk tolerance level, $\zeta$ as input. The lower a user’s risk tolerance, the more the user wants to avoid the cost arising from server failure. The higher a user’s risk tolerance, the less the user is willing to prepare for failure, as it incurs an additional storage and migration cost. Next, this algorithm converts $Q_{riskAware}$ to a vector of probabilities $p_{riskAware}$, where $Q_{riskAware}$ is the converged policy trained from one of our importance-sampling failure-adaptive Q-learning simulator algorithms (Algorithms 1, 3, 4 or 2). The algorithm also converts $Q_{riskTaking}$ to a vector of probabilities $p_{riskTaking}$, where $Q_{riskTaking}$ is the converged policy of a Q-learning algorithm that does not take into account rare events in modelling of their environment, and does not consider the possibility of backups. This represents the risk taking component, for users who do not want to plan for rare events during service migration.
Algorithm 5 is an online algorithm to be used in real-time scenarios. For each timeslot, the state transition occurs according to $p(s^{\prime}|s,a)$, and rare events according to the true rare events probability $\epsilon(s)$. At each timeslot, the policy which is used will be in accordance with the risk tolerance level $\zeta\in[0,1]$. The lower the risk tolerance $\zeta$, the higher the probability that the action selected will be sampled according to $p_{riskAware}$ (lines 6-10), which is obtained from our algorithm’s converged policy. When $\zeta=0$, the action selected will be according to our algorithm’s converged policy. In the following theorem, we characterize the condition under which it would have been beneficial to prepare for rare events.
@@
After $T$ time-slots, at risk level $\zeta$, and under our algorithm, assuming our algorithm is able to prevent the cost of failures with probability $1-f$, it benefited the user to prepare for rare events, if
@@
(a) Convergence of tabular importance sampling Q-learning algorithm ImRE.
(b) Convergence of actor critic importance sampling algorithm ImACRE
(c) Convergence of deep Q-learning importance sampling algorithm ImDQL.
Figure 3: Convergence of our proposed importance-sampling based rare events adaptive algorithms. Here, ImRE and ImACRE are applied to the a special case of the scenario (Section 3) where every user has their own service profile (the single user case), and ImDQL is applied to the scenario where users share service profiles (Section 4). All three variations of our algorithm converge.
@@
Figure 4: Single user service migration scenario: Comparison of our actor critic algorithm importance sampling ImACRE, with actor critic versions of the baselines NIS, WBA and RES, in an online scenario.ImACRE leads to lower costs on average and in rare failure states, but higher storage and delay costs in normal states.

### v3

@@ 6.3 Boundedness and Convergence Properties
See our Technical Report [32].
The main idea of the proof is as follows. Firstly, we show that the optimal policy of the true system is a fixed point of the following equation with our simulator’s transition probability $\hat{p}(s^{\prime}|s,a)$ and the importance sampling correction weight $w_{t}(s)$: $Q^{*}(s,a)=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}Q^{*}(s^{\prime},b)].$ We define the operator $\mathbf{H}\hat{Q}=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}(s^{\prime},b)]$, and having shown that that $\mathbf{H}Q^{*}=Q^{*}$, we next show that $\mathbf{H}\hat{Q}$ is a contraction mapping, through showing that $||\mathbf{H}\hat{Q}_{1}-\mathbf{H}\hat{Q}_{2}||_{\infty}\leq\gamma||\hat{Q}_{1}-\hat{Q}_{2}||_{\infty}$. Next, we invoke the result of Corollary 3 on the convergence of random processes. We rewrite the update equation of Algorithm 1, such that it fits the form of the equations in Corollary 3, and show that the conditions of Corollary 3 hold. We do this by letting $G_{t}(s,a)=w^{t}[r(s,a,X(s,a))+\gamma\max_{b}\hat{Q}^{t}(X(s,a),b)]-Q^{*}(s,a)$, and by using the fact that $\mathbf{H}\hat{Q}$ is a contraction mapping, we show that $||\mathbb{E}[G_{t}(s,a)]||_{\infty}\leq\gamma||\hat{Q}^{t}(s,a)-Q^{*}||_{\infty}=\gamma||\triangle_{t}||_{\infty}$, and $\mathbf{var}[G_{t}(s,a)]\leq C(1+||\triangle_{t}||^{2}_{\infty})$.
@@
Rare Event Adaptive Eligibility Traces Algorithm (FIRE-ETAA): To speed up convergence for the tabular Q learning algorithm, we propose another variant of Algorithm 1 using eligibility traces. This algorithm (ETAA) is a combination of our proposed importance sampling based Q-learning (Algorithm 1) and Watkin’s $Q(\lambda)$ [35], which is an algorithm which combines Q-learning with eligibility traces. We present this algorithm in our Technical Report [32].
7 Scaling to Large State or Action Spaces
@@
To handle large state spaces, as might be found in real-world edge computing networks that span one or more cities and surrounding suburbs, we propose two variants of Algorithm 1 using deep Q-learning (Algorithm FIRE-ImDQL) and actor critic (Algorithm FIRE-ImACRE) respectively. The codes are available online at [36].
@@
Rare Event Adaptive Deep Q-learning algorithm (FIRE-ImDQL, Algorithm 2): We modify and combine the classic deep reinforcement learning algorithm in [37] with our importance sampling based rare events adaptive algorithm (Algorithm 1) by proposing an importance sampling based loss function.
@@
Algorithm 2 FIRE-ImDQL: Importance Sampling for Rare Events using Deep Q Learning
@@
$L(\boldsymbol{\theta})=\mathbb{E}_{s,a,s^{\prime}}[w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))-Q(s,a,\boldsymbol{\theta})]^{2},$
where $\boldsymbol{\theta}^{-}$ are the weights of the target neural network, $\boldsymbol{\theta}$ are the weights of the predicted neural network, which are updated every iteration, and $w_{t}$ is the importance sampling weight. The loss function is optimized via gradient descent in line 13. We are trying to minimize the difference between the target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ and the prediction $Q(s,a,\boldsymbol{\theta})$. The target network is updated only every $K$ iterations (line 16), while the prediction network is updated every iteration. Having the target fixed for awhile helps in having stable training. The key idea which differentiates our algorithm from traditional deep reinforcement learning algorithms is that our target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ is “error corrected” by the importance sampling weight $w_{t}$, and the rare event distribution is updated via $T,U$ and $\hat{\epsilon}$. The variables $T$, $U$, $\hat{\epsilon}$ and $w_{t}$ are updated in the same manner as the tabular algorithm (Algorithm 1), through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t+1}$. The actions are also selected in the same manner as Algorithm 1, through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t}$.
Algorithm 3 FIRE-ImACRE: Importance Sampling based Advantage Actor Critic for Rare Events
@@
Rare Event Adaptive Actor Critic Algorithm (FIRE-ImACRE, Algorithm 3): Neural networks are used to approximate the “actor” and “critic” functions [38]. The critic function estimates the value (network $c$), based on which the actor function updates the policy (network $\mathbf{\theta}$). The advantage function $A(s^{t},a^{t})$ (line 8) measures how good taking action $a^{t}$ is, compared to taking the average action, at state $s^{t}$. It is a function of the reward $r^{t+1}$, the values $V(s^{t})$ and $V(s^{t+1})$ (obtained from the critic network $\boldsymbol{c}$ , line 7). It is corrected by the importance sampling correction weight $w^{t}(s^{t},a^{t},s^{t+1})$ (line 8), because we perform importance sampling of rare events. Each iteration, $A(s^{t},a^{t})$ will be stored in the buffer (line 9). The variables $T,U,\hat{\epsilon}$ and $w_{t}$ are updated in the same manner as the tabular algorithm (Algorithm 1). At the end of an epoch ($K$ timeslots), both the actor and critic networks are optimized, via gradient descent. The gradients for the actor and critic networks, respectively, are:
@@
8 Heterogeneous Risk Tolerances
Edge computing users, application vendors, and network providers have different tolerances towards server failures and the resulting higher latency which occur. Some applications are highly latency sensitive, and any server failure has a significant impact on safety and smooth functioning of the application, e.g., obstacle detection in autonomous vehicles, or immersive outdoor sport applications. Other users with less urgent jobs may not priortize a failure-aware backup placement solution, due to the extra costs of storing and migrating backups. To deal with heterogeneous risk tolerances, one potential solution is to alter the failure cost when training the algorithm. Nevertheless, the same rare event with the same system cost may lead to varying risk tolerances and willingness to incur a preparation cost amongst different users. Therefore, we propose a mechanism that can cater to edge computing users of heterogeneous risk tolerances. It takes the users’ risk tolerance towards server failures as an input, and derives a joint service placement and backup placement policy, based on the users’ risk tolerance. With RiTA (Risk Tolerance-Adaptive Algorithm), we also do not need to re-train a new policy $Q_{riskAware}$ for every possible level of risk, which scales better to a wider variety of users.
@@ 8 Heterogeneous Risk Tolerances
We present this algorithm in Algorithm 4 (RiTA). Firstly, RiTA takes the user’s risk tolerance level, $\zeta$ as input. The lower a user’s risk tolerance, the more the user wants to avoid the cost arising from server failure. The higher a user’s risk tolerance, the less the user is willing to prepare for failure. Next, this algorithm converts $Q_{riskAware}$ to a vector of probabilities $p_{riskAware}$, where $Q_{riskAware}$ is the converged policy trained from one of our importance-sampling failure-adaptive Q-learning simulator algorithms (Algorithms 1, 2, or 3). The algorithm also converts $Q_{riskTaking}$ to a vector of probabilities $p_{riskTaking}$, where $Q_{riskTaking}$ is the converged policy of a Q-learning algorithm that does not take into account rare events in modelling of their environment, and does not consider the possibility of backups. This represents the risk taking component. Algorithm 4 is an online algorithm for real-time scenarios. For each timeslot, the state transition occurs according to $p(s^{\prime}|s,a)$, and rare events according to the true rare events probability $\epsilon(s)$. At each timeslot, the policy which is used will be in accordance with the risk tolerance level $\zeta\in[0,1]$. The lower the risk tolerance $\zeta$, the higher the probability that the action selected will be sampled according to $p_{riskAware}$ (lines 6-10), obtained from our algorithm’s converged policy. When $\zeta=0$, the action selected will be according to our algorithm’s converged policy. In the next theorem, we characterize the condition under which it would have been beneficial to prepare for rare events.
@@
After $T$ time-slots, at risk level $\zeta$, assuming our algorithm is able to prevent the cost of failures with probability $1-f$, it benefited the user to prepare for rare events, if
@@
(a) Convergence of tabular importance sampling Q-learning algorithm FIRE-ImRE.
(b) Convergence of actor critic importance sampling algorithm FIRE-ImACRE
(c) Convergence of deep Q-learning importance sampling algorithm FIRE-ImDQL.
Figure 3: Convergence of our proposed importance-sampling based rare events adaptive algorithms. Here, FIRE-ImRE and FIRE-ImACRE are applied to the a special case of the scenario (Section 3) where every user has their own service profile (the single user case), and FIRE-ImDQL is applied to the scenario where users share service profiles (Section 4). All three variations of our algorithm converge.
@@
Figure 4: Single user service migration scenario: Comparison of our actor critic algorithm importance sampling FIRE-ImACRE, with actor critic versions of the baselines NIS, WBA and RES, in an online scenario.FIRE-ImACRE leads to lower costs on average and in rare failure states, but higher storage and delay costs in normal states.